# Reasoning is All You Need for Video Generalization: A Counterfactual Benchmark with Sub-question Evaluation

### v1

@@ Abstract
Counterfactual reasoning is crucial for robust video understanding but remains underexplored in existing multimodal benchmarks. In this paper, we introduce COVER (COunterfactual VidEo Reasoning), a multidimensional multimodal benchmark that systematically evaluates MLLMs across the abstract-concrete and perception-cognition dimensions. Beyond prior multimodal benchmarks, COVER decomposes complex queries into structured sub-questions, enabling fine-grained reasoning analysis. Experiments on commercial and open-source models reveal a strong correlation between sub-question accuracy and counterfactual reasoning performance, highlighting the role of structured inference in video understanding. Furthermore, our results suggest a key insight: enhancing the reasoning capability of models is essential for improving the robustness of video understanding. COVER establishes a new standard for assessing MLLMs’ logical reasoning abilities in dynamic environments.
@@
Qiji Zhou11footnotemark: 1 1, YiFan Gong11footnotemark: 1 2, Guangsheng Bao1, Hongjie Qiu2, Jinqiang Li2 Xiangrong Zhu2, Huajian Zhang1, Yue Zhang22footnotemark: 2 1 1School of Engineering, Westlake University 2College of Computer Science and Technology, Hangzhou Dianzi University {zhouqiji, baoguangsheng, zhanghuajian, zhangyue}@westlake.edu.cn {gongyifan, qiuhongjie, lijinqiang, zhuxiangrong}@hdu.edu.cn
@@ 1 Introduction
In recent years, the rapid advancement of large language models (LLMs) has spurred growing interest in multimodal large language models (MLLMs) Hurst et al. (2024); Anthropic (2024); Chen et al. (2025); Zhang et al. (2024, 2025); Wang et al. (2024); Wu et al. (2024b). Various early benchmarks have been proposed to assess multimodal understanding ability of MLLMs, particularly in static images Fu et al. (2024a); Hudson and Manning (2019); Liu et al. (2024); Yu et al. (2024). More recently, benchmarks involving complex images and dynamic videos have emerged to evaluate MLLM’s capabilities in temporal reasoning, spatio-temporal recognition, and object detection Fu et al. (2024b); Li et al. (2024b, 2023). Despite these advances, these benchmarks often overlook counterfactual reasoning, which is a critical component for evaluating inference in complex and realistic environments. As a result, they fall short of providing a comprehensive assessment of MLLMs’ reasoning capabilities.
@@ 2 Related Work
MLLMs and Their Evaluation. Recent advancements in Multimodal Large Language Models (MLLMs) have significantly enhanced their ability to process multimodal data, such as text and images, or text and video. Specialized evaluation benchmarks have been developed to assess models’ understanding of static images, such as MM-Vet Yu et al. (2024), MME Fu et al. (2024a), MMBench Liu et al. (2024), and GQA Hudson and Manning (2019), which primarily focus on visual comprehension. Extending these to dynamic multimodal reasoning, video-based benchmarks like Video-MME Fu et al. (2024b), Mvbench Li et al. (2024b), and SEED-Bench Li et al. (2023) evaluate models’ temporal and contextual understanding. These benchmarks play a crucial role in evaluating the capabilities of MLLMs in dynamic and evolving environments, highlighting the growing emphasis on multimodal reasoning across diverse tasks.
Chain-of-Thought and Counterfactual Reasoning in MLLMs. Chain-of-Thought (CoT) reasoning Wei et al. (2022) decomposes complex problems into sequential steps, improving logical reasoning. In MLLMs, multimodal CoT variants Zhang et al. (2023b); Zheng et al. (2023) have shown notable improvements in reasoning across modalities. Furthermore, counterfactual reasoning, which investigates how modifying specific conditions affects outcomes, has gained increasing attention. Previous studies have evaluated LLMs on text-based counterfactual tasks Wu et al. (2024c, a) and visual counterfactual Q&A Li et al. (2024e). ACQUIRED Wu et al. (2023) offers a comprehensive taxonomy for counterfactual scenarios, enriching this area of research. Additionally, CoFCA Wu et al. (2024a) incorporates sub-problem evaluation to assess evidence integration and multi-step reasoning capabilities. CoT and counterfactual reasoning provide a more systematic and effective approach to multimodal reasoning, significantly enhancing the models’ ability to handle complex reasoning tasks.
Generalization of Multimodal Reasoning and Video Counterfactual Tasks. Despite significant contributions from video counterfactual benchmarks like CRIPP-VQA, which focuses on physical attributes, VITATECS, which emphasizes captioning, and ACQUIRED’s limited counterfactual taxonomy Li et al. (2024c); Patel et al. (2022), their scope remains limited.
To address these gaps, COVER introduces sub-question decomposition, allowing the evaluation of perceptual and cognitive reasoning in counterfactual video tasks. As shown in Table 1, COVER spans four dimensions—distinguishing between abstract and concrete object attributes and perceptual and cognitive demands offering a more comprehensive evaluation framework than existing benchmarks.
@@ 3 The COVER Benchmark
This section provides a comprehensive overview of the construction of COVER. We introduce its data partitioning framework designed to evaluate MLLM reasoning ability across four complementary dimensions. Next, we describe the data curation process, which domain experts have rigorously validated to ensure the high quality and reliability of the benchmark.
@@ 3.2 Data Construction
To construct COVER, we carefully selected a diverse range of open-source and research-available video sources, including but not limited to Sigurdsson et al. (2016); Yi et al. (2020); Xie et al. (2024); Tan et al. (2020); Shahroudy et al. (2016); Patraucean et al. (2024); Zhang et al. (2023a); Gao et al. (2017); Jang et al. (2017); Wang et al. (2019); Krantz et al. (2020). These sources encompass various real-world scenarios, ranging from daily activity recognition to complex scene understanding. As shown in Appendix Figure 7, we collected 146 videos and designed 150 aspect-specific QA pairs, each of which underwent dual-team review for validation. To ensure balanced coverage across the four quadrants, we expanded the seed data using GPT-generated instances (720-760 per quadrant) to mitigate any potential biases. The detailed statistical findings are comprehensively presented in Figure 3. The frame count of videos in COVER ranges from 16 to 1739, with an average of 294.34 frames. We finally constructed 2,923 high-quality counterfactual question-answer pairs. Each question-answer pair consists of an original question, which presents no hypothetical context, and a counterfactual question, which introduces situational assumptions and sub-questions that enable granular reasoning analysis.
Eight annotators further validated the dataset and checked logical consistency to ensure the reasoning relied solely on the video content. Additionally, three experts cross-valid the dataset (see Appendix Table 6) to confirm the structural balance.
@@ 4 Experiments
In this section, we conduct a thorough evaluation of MLLMs of varying scales on the COVER dataset. We assess model performance across four key quadrants, ranging from perception to cognition and from abstract to concrete. Then, we evaluate their capabilities in counterfactual reasoning, direct reasoning, and sub-question-guided reasoning across multiple subtasks. The results include a detailed comparison of open-source and commercial closed-source models of different sizes.
@@ 4.1 Settings
To ensure a comprehensive and in-depth evaluation, we selected a diverse set of representative MLLMs, including both commercial closed-source models such as GPT-4o Hurst et al. (2024), Claude Anthropic (2024), and Gemini Team et al. (2024), as well as prominent open-source models like InternVL2.5 Chen et al. (2025), LLAVA-Video Zhang et al. (2024), LLaVA-OV Li et al. (2024a), Qwen2-VL Wang et al. (2024), VideoLLaMA3 Zhang et al. (2025), vila-u Wu et al. (2024b), and VILA Lin et al. (2023). These models represent a broad spectrum of parameter scales and technical characteristics, providing a well-rounded reflection of technological development in multimodal learning. We use , , and to evaluate the original question accuracy, counterfactual question accuracy, and sub-question accuracy of MLLM in video understanding, using the average score from at least three runs. We adopted a consistent frame extraction strategy for the video inputs, sampling 16 frames per segment, ensuring that all models are evaluated under identical conditions. In addition, we will explore the effect of different sampling strategies in Chapter 5.
Models AC (%) CC (%) CP (%) AP (%) GPT-4o 71.05 41.81 41.70 74.87 43.65 68.36 69.95 42.62 50.52 65.01 55.65 63.97 GPT-4o-mini 62.29 52.40 42.97 76.32 54.37 66.49 64.62 40.85 44.78 65.56 58.26 65.96 Claude-3.5-sonnet 56.92 37.01 35.55 70.11 42.33 61.77 60.03 32.88 40.08 66.94 39.81 56.81 Gemini 1.5 Pro 69.49 44.49 53.36 82.14 51.98 72.78 71.76 43.93 56.81 75.48 57.99 69.54 Gemini 1.5 Flash 70.48 45.34 52.23 82.01 49.34 71.51 70.67 42.02 51.90 72.04 58.26 71.36 Gemini 2.0 Flash 74.29 44.36 51.38 83.99 47.75 72.84 74.22 38.74 58.26 75.90 57.71 66.84 InternVL2.5-78B 72.88 59.60 57.67 80.95 63.62 75.62 75.99 58.25 63.65 76.86 56.20 70.07 LLaVA-Video-72B 53.11 54.94 53.14 65.34 60.45 67.03 67.94 52.39 53.49 70.66 56.20 70.01 InternVL2.5-26B 71.05 47.74 50.53 80.95 58.99 72.17 76.13 47.20 60.12 73.14 50.00 65.61 InternVL2.5-8B 69.77 58.62 49.96 80.95 64.55 71.02 73.94 55.80 54.66 72.18 51.79 68.19 VideoLLama3-8B 68.08 45.62 49.68 81.35 54.89 68.36 72.99 50.75 51.62 69.28 53.44 67.90 LLaVA-ov-7B 54.66 51.69 47.49 62.96 53.04 61.77 64.53 49.66 49.48 68.60 52.75 64.73 LLaVA-Video-7B 50.14 55.23 44.52 61.64 50.53 60.50 63.57 52.52 49.97 66.39 49.59 63.03 Qwen2-VL-7B 65.96 49.15 48.41 82.14 43.39 67.03 71.21 45.57 50.52 67.49 49.72 65.02 VILA-U-7B 58.19 39.83 38.87 63.10 41.93 54.51 59.07 37.93 37.94 59.50 33.88 55.34 VILA1.5-7B 54.80 55.93 39.29 66.93 62.30 63.52 55.25 58.53 44.64 63.64 52.34 61.91
@@ 4.2 Main Results
Open-source Model Performance. As shown in Table 3, InternVL2.5-78B leads in abstract cognition ( 72.88%) and concrete perception tasks ( 58.25%), reflecting a deep understanding of abstract concepts and complex logic. Lightweight models like Qwen2-VL-7B perform well in concrete cognition ( 82.14%) but face limitations in abstract tasks ( 65.96% in AC) due to their smaller parameter size, revealing distinct capabilities across model types. Commercial models, such as the Gemini series, maintain strong performance in concrete cognition and abstract perception tasks but generally fall behind open-source models in counterfactual reasoning. Most models struggle with counterfactual reasoning, with only InternVL2.5-7BB and VILA1.5-7B showing some task-specific advantages, highlighting the need for targeted optimization in conditional hypothesis modeling.
@@ 4.3 Sub-question Guideline
The column in Guide-CoT represents sub-questions containing standard answers. For InternVL2.5-78B, reaches 68.29%, showing an 8.63% improvement over the no-CoT baseline, compared to CoT’s mere 0.96% enhancement (59.46% → 60.42%). This demonstrates that providing complete answers significantly enhances reasoning accuracy, especially in complex or multi-step tasks. Sub-questions with standard answers help the model better integrate information and validate reasoning steps, leading to improved consistency and performance. In Appendix Figure 9, we provide detailed case demonstrations to further explore these observations and examine the relationships between reasoning paths and task complexities.
@@ 5.1 Ablation Study of Video Frames
5.2 An Analysis of Robustness and Logical Reasoning Capability in MLLMs
@@
The ability of MLLMs to answer original questions serves as a key indicator of their overall understanding capabilities, while performance on sub-questions reveals single-step reasoning proficiency. A notable observation is the strong Pearson correlation between and reaches 0.836, indicating a strong connection between model understanding and reasoning capabilities. Furthermore, as shown in Table 2, the correlation between and is 0.608. These moderately strong correlations indicate that a model’s ability to comprehend original questions plays a fundamental role in enabling effective step-by-step reasoning. similarly, the correlation between and suggests that models with a higher understanding capability tend to perform better when solving decomposed sub-questions, reinforcing the notion that comprehension and reasoning are interdependent. However, the moderate correlation between and suggests that counterfactual reasoning involves additional complexities, making it a more challenging task than single-step reasoning.
@@
Heatmaps in Figure 4 reveal distinct performance patterns across the different quadrants, and the relationship between comprehension, step-by-step reasoning, and counterfactual inference. In the abstract cognition quadrant, which includes high-level reasoning tasks, the performance drop from to is minimal, and the transition to remains stable. This suggests that while sub-question reasoning is important, models still maintain strong performance for abstract reasoning tasks even when counterfactual assumptions are introduced. Conversely, in the concrete perception quadrant, a more pronounced performance decline is observed from to , and even further to . This indicates that sub-question reasoning has a more significant impact on counterfactual reasoning in perception-based tasks, where models struggle to decompose complex perceptual information into manageable steps.
@@
Overall, these findings suggest that counterfactual reasoning is inherently more complex than single-step reasoning, particularly in perception-intensive tasks where models must infer causal relationships rather than relying on pattern recognition alone. The relatively stable transition between and in abstract cognition tasks implies that models leverage their existing conceptual understanding to handle counterfactual scenarios more effectively. Improving counterfactual reasoning in perception-heavy tasks remains a key challenge, likely requiring enhancements in reasoning mechanisms and causal inference capabilities.

### v2

@@ Abstract
Counterfactual reasoning is crucial for robust video understanding but remains underexplored in existing multimodal benchmarks. In this paper, we introduce COVER (COunterfactual VidEo Reasoning), a multidimensional multimodal benchmark that systematically evaluates MLLMs across the abstract-concrete and perception-cognition dimensions. Beyond prior multimodal benchmarks, COVER decomposes complex queries into structured sub-questions, enabling fine-grained reasoning analysis. Experiments on commercial and open-source models reveal a strong correlation between sub-question accuracy and counterfactual reasoning performance, highlighting the role of structured inference in video understanding. Furthermore, our results suggest a key insight: enhancing the reasoning capability of models is essential for improving the robustness of video understanding. COVER establishes a new standard for assessing MLLMs’ logical reasoning abilities in dynamic environments. Our work is available at https://github.com/gongyifan-hash/COVER-Benchmark.
@@
Qiji Zhou111Equal contribution. 1, YiFan Gong111Equal contribution. 2, Guangsheng Bao1, Hongjie Qiu2, Jinqiang Li2 Xiangrong Zhu2, Huajian Zhang1, Yue Zhang222Corresponding Author. 1 1School of Engineering, Westlake University 2College of Computer Science and Technology, Hangzhou Dianzi University {zhouqiji, baoguangsheng, zhanghuajian, zhangyue}@westlake.edu.cn {gongyifan, qiuhongjie, lijinqiang, zhuxiangrong}@hdu.edu.cn
@@ 1 Introduction
In recent years, the rapid advancement of large language models (LLMs) has spurred growing interest in multimodal large language models (MLLMs) Hurst et al. (2024); Anthropic (2024); Chen et al. (2024); Zhang et al. (2024a, 2025); Wang et al. (2024); Wu et al. (2024b). Various early benchmarks have been proposed to assess multimodal understanding ability of MLLMs, particularly in static images Fu et al. (2023); Hudson and Manning (2019); Liu et al. (2024); Yu et al. (2024). More recently, benchmarks involving complex images and dynamic videos have emerged to evaluate MLLM’s capabilities in temporal reasoning, spatio-temporal recognition, and object detection Fu et al. (2024); Li et al. (2024b, 2023). Despite these advances, these benchmarks often overlook counterfactual reasoning, which is a critical component for evaluating inference in complex and realistic environments. As a result, they fall short of providing a comprehensive assessment of MLLMs’ reasoning capabilities.
@@ 2 Related Work
Multimodal Large Language Models and Their Evaluation. Recent advances in MLLMs have greatly improved their capacity to understand and reason over diverse modalities, such as images, text, and videos. To evaluate these models, benchmarks targeting static image comprehension have emerged, including MM-Vet Yu et al. (2024), MME Fu et al. (2023), MMBench Liu et al. (2024), and GQA Hudson and Manning (2019). These primarily assess visual recognition and spatial reasoning. Extending beyond static content, video-centric benchmarks like Video-MME Fu et al. (2024), MvBench Li et al. (2024b), and SEED-Bench Li et al. (2023) focus on temporal dynamics and contextual reasoning. Together, these benchmarks reflect the growing demand for evaluating multimodal understanding in both static and dynamic environments.
Chain-of-Thought and Counterfactual Reasoning in MLLMs. Chain-of-Thought (CoT) reasoning Wei et al. (2022) enhances logical inference by breaking down complex tasks into intermediate steps. Multimodal adaptations Zhang et al. (2024b); Zheng et al. (2023) extend this strategy across modalities, showing gains in structured reasoning. Counterfactual reasoning, which examines hypothetical changes and their consequences, has also gained traction. Prior work explores this in text Wu et al. (2024c, a), visual QA Li et al. (2024e), and hybrid settings. ACQUIRED Wu et al. (2023) proposes a taxonomy of counterfactual types, while AuroraCap Chai et al. (2024) and CoFCA Wu et al. (2024a) assess models’ sub-task decomposition and multi-step reasoning. These approaches collectively underscore the importance of structured, causal reasoning in complex multimodal tasks.
Multimodal Generalization and Video Counterfactual Benchmarks. Although several benchmarks target video-based counterfactual understanding—such as CRIPP-VQA for physical properties, VITATECS for captioning, and ACQUIRED for scenario taxonomy Li et al. (2024c); Patel et al. (2022)—they remain narrow in scope. Most fail to capture the breadth of reasoning demands in real-world counterfactual scenarios.
To address this, COVER introduces a fine-grained framework for evaluating counterfactual video reasoning via sub-question decomposition. It explicitly distinguishes between abstract vs. concrete object attributes and perceptual vs. cognitive reasoning demands. As summarized in Table 1, COVER broadens the evaluation spectrum, enabling a more nuanced and comprehensive assessment of multimodal counterfactual reasoning than prior efforts.
@@ 3 The COVER Benchmark
This section provides a comprehensive overview of the construction of COVER. We introduce our data partitioning framework designed to evaluate MLLM reasoning ability across four complementary dimensions. Next, we describe the data curation process, which domain experts have rigorously validated to ensure the high quality and reliability of the benchmark.
Our benchmark includes approximately 2,800 videos, which are paired with around 12,000 to 13,000 individual QA instances. As stated in L-Figure 6, the enhanced version of our dataset consists of about 2.9k question pairs, with each pair comprising at least three individual QA items: one original question, one counterfactual question, and at least one sub-question (often multiple).
@@ 3.2 Data Construction
To construct COVER, we carefully selected a diverse range of open-source and research-available video sources, including Sigurdsson et al. (2016); Yi et al. (2020); Xie et al. (2024); Tan et al. (2020); Shahroudy et al. (2016); Pătrăucean et al. (2023); Zhang et al. (2023); Gao et al. (2017); Jang et al. (2017); Wang et al. (2019); Krantz et al. (2020). These sources encompass various real-world scenarios, ranging from daily activity recognition to complex scene understanding. As shown in Appendix Figure 6, we collected 146 videos and designed 150 aspect-specific QA pairs, each of which underwent dual-team review for validation. To ensure balanced coverage across the four quadrants, we expanded the seed data using GPT-generated instances (720-760 per quadrant) to mitigate any potential biases. The detailed statistical findings are comprehensively presented in Figure 3. The frame count of videos in COVER ranges from 16 to 1739, with an average of 294.34 frames. We finally constructed 2,923 high-quality counterfactual question-answer pairs. Each question-answer pair consists of an original question, which presents no hypothetical context, and a counterfactual question, which introduces situational assumptions and sub-questions that enable granular reasoning analysis.
Eight annotators further validated the dataset and checked logical consistency to ensure the reasoning relied solely on the video content. Additionally, three experts cross-validate the dataset (see Appendix Table 9) to confirm the structural balance.
@@ 4 Experiments
In this section, we systematically evaluate MLLMs of varying scales on the COVER benchmark to foster transparent and reproducible research. Our evaluation spans four key dimensions: cognition, perception, abstraction, and concreteness. It encompasses diverse reasoning sub-tasks, including counterfactual reasoning, direct inference, and sub-question-guided reasoning. We compare both open-source and proprietary models across different parameter scales to analyze their relative strengths and limitations. We begin by detailing the experimental setup.
@@ 4.1 Settings
To ensure a thorough evaluation, we selected a diverse set of representative MLLMs, including commercial closed-source models such as GPT-4o Hurst et al. (2024), Claude Anthropic (2024), and Gemini Reid et al. (2024), as well as leading open-source models such as InternVL2.5 Chen et al. (2024), LLAVA-Video Zhang et al. (2024a), LLaVA-OV Li et al. (2024a), Qwen2-VL Wang et al. (2024), VideoLLaMA3 Zhang et al. (2025), vila-u Wu et al. (2024b), and VILA Lin et al. (2024). These models span a wide range of parameter scales and design paradigms, offering a comprehensive view of the current landscape in multimodal learning.
We evaluate model performance on video understanding using three metrics: (original question accuracy), (counterfactual question accuracy), and (sub-question accuracy), with scores averaged over at least three runs. All models are tested under identical conditions, using a consistent frame extraction strategy that samples 16 frames per video segment. The impact of alternative sampling strategies is discussed in Chapter 5.
Models A&C (%) C&C (%) C&P (%) A&P (%) GPT-4o 71.05 41.81 41.70 74.87 43.65 68.36 69.95 42.62 50.52 65.01 55.65 63.97 GPT-4o-mini 62.29 52.40 42.97 76.32 54.37 66.49 64.62 40.85 44.78 65.56 58.26 65.96 Claude-3.5-sonnet 56.92 37.01 35.55 70.11 42.33 61.77 60.03 32.88 40.08 66.94 39.81 56.81 Gemini 1.5 Pro 69.49 44.49 53.36 82.14 51.98 72.78 71.76 43.93 56.81 75.48 57.99 69.54 Gemini 1.5 Flash 70.48 45.34 52.23 82.01 49.34 71.51 70.67 42.02 51.90 72.04 58.26 71.36 Gemini 2.0 Flash 74.29 44.36 51.38 83.99 47.75 72.84 74.22 38.74 58.26 75.90 57.71 66.84 InternVL2.5-78B 72.88 59.60 57.67 80.95 63.62 75.62 75.99 58.25 63.65 76.86 56.20 70.07 LLaVA-Video-72B 53.11 54.94 53.14 65.34 60.45 67.03 67.94 52.39 53.49 70.66 56.20 70.01 InternVL2.5-26B 71.05 47.74 50.53 80.95 58.99 72.17 76.13 47.20 60.12 73.14 50.00 65.61 InternVL2.5-8B 69.77 58.62 49.96 80.95 64.55 71.02 73.94 55.80 54.66 72.18 51.79 68.19 VideoLLama3-8B 68.08 45.62 49.68 81.35 54.89 68.36 72.99 50.75 51.62 69.28 53.44 67.90 LLaVA-ov-7B 54.66 51.69 47.49 62.96 53.04 61.77 64.53 49.66 49.48 68.60 52.75 64.73 LLaVA-Video-7B 50.14 55.23 44.52 61.64 50.53 60.50 63.57 52.52 49.97 66.39 49.59 63.03 Qwen2-VL-7B 65.96 49.15 48.41 82.14 43.39 67.03 71.21 45.57 50.52 67.49 49.72 65.02 VILA-U-7B 58.19 39.83 38.87 63.10 41.93 54.51 59.07 37.93 37.94 59.50 33.88 55.34 VILA1.5-7B 54.80 55.93 39.29 66.93 62.30 63.52 55.25 58.53 44.64 63.64 52.34 61.91
@@ 4.2 Main Results
Open-source Model Performance. As shown in Table 3, InternVL2.5-78B leads in abstract cognition ( 72.88%) and concrete perception tasks ( 58.25%), reflecting a deep understanding of abstract concepts and complex logic. Lightweight models like Qwen2-VL-7B perform well in concrete cognition ( 82.14%) but face limitations in abstract tasks ( 65.96% in A&C) due to their smaller parameter size, revealing distinct capabilities across model types. Commercial models, such as the Gemini series, maintain strong performance in concrete cognition and abstract perception tasks but generally fall behind open-source models in counterfactual reasoning. Most models struggle with counterfactual reasoning, with only InternVL2.5-7BB and VILA1.5-7B showing some task-specific advantages, highlighting the need for targeted optimization in conditional hypothesis modeling.
@@ 4.3 Sub-question Guideline
The column in Guide-CoT indicates sub-questions that include standard answers. For InternVL2.5-78B, reaches 68.29%, reflecting an 8.63% improvement over the no-CoT baseline, in contrast to CoT’s modest gain of only 0.96% (from 59.46% to 60.42%). This suggests that providing complete answers substantially enhances reasoning accuracy, particularly in complex or multi-step tasks. Standard-answer sub-questions enable the model to better integrate information and verify intermediate reasoning steps, resulting in improved consistency and overall performance. Detailed case studies are presented in Appendix Figure 9 to further illustrate these findings and analyze the interplay between reasoning paths and task complexity.
@@ 5.1 Ablation Study of Video Frames
5.2 Robustness and Logical Reasoning in MLLMs
@@
The ability of MLLMs to answer original questions serves as a key indicator of their overall understanding capabilities, while performance on sub-questions reveals single-step reasoning proficiency. A notable observation is the strong Pearson correlation between and reaches 0.836, indicating a strong connection between model understanding and reasoning capabilities. Furthermore, as shown in Figure 5, the correlation between and is 0.608. These moderately strong correlations indicate that a model’s ability to comprehend original questions plays a fundamental role in enabling effective step-by-step reasoning. Similarly, the correlation between and suggests that models with a higher understanding capability tend to perform better when solving decomposed sub-questions, reinforcing the notion that comprehension and reasoning are interdependent. However, the moderate correlation between and suggests that counterfactual reasoning involves additional complexities, making it a more challenging task than single-step reasoning.
As illustrated in Table 6, We observed that across multiple models, the probability was significantly higher than , clearly indicating that the correctness of sub-questions is a strong predictor of overall counterfactual performance.
Model | P(|) | P(|) | P(|) | P(|) |
gemini-1.5-pro | 56.54 | 43.45 | 44.99 | 55.01 |
GPT-4o-mini | 59.49 | 40.51 | 47.65 | 52.35 |
InternVL2.5-78B | 62.90 | 37.10 | 56.67 | 43.34 |
LlaVA-Video-72B | 63.28 | 36.72 | 51.60 | 48.40 |
@@
Analysis of the heat maps in Figure 4 reveals different performance patterns in the quadrants, highlighting the interaction between comprehension, step-by-step reasoning, and counterfactual inference. In abstract reasoning tasks such as social inference and procedural understanding, the drop from to is minimal, and the transition to remains stable. This suggests that models can effectively leverage sub-question reasoning and maintain performance even under counterfactual assumptions. In contrast, the concrete perception quadrant—involving tasks like object recognition and motion understanding—shows a sharper decline from to , and further to . This indicates that perception-heavy tasks pose greater challenges, as models struggle to decompose complex sensory input into reasoning steps required for counterfactual understanding.
@@
Overall, our findings indicate that counterfactual reasoning is inherently more challenging than single-step reasoning, especially in perception-intensive tasks where models must infer causality beyond pattern recognition. In contrast, the relatively stable gap between and in abstract-cognitive tasks suggests that models can better leverage conceptual knowledge. Enhancing counterfactual reasoning in perception-heavy scenarios remains a key challenge, likely requiring improved causal inference and reasoning mechanisms.
5.3 The Effects of Model Scale
We conduct systematic analyses to characterize performance gaps across original, counterfactual, and sub-question accuracies. Our goal is to mitigate these gaps by examining factors such as model scale, training alignment, and reasoning strategies. As shown in Table 7, with similar visual backbones, increasing language model size significantly reduces the performance gap—particularly between sub-question and counterfactual accuracy. Specifically, the absolute difference between (70.07%) and (40.68%) is 29.39% for the 2B model, increases slightly to 30.38% for the 4B model, and then drops substantially to 16.56% for the 8B model. Similarly, the gap between and grows from 14.81% (2B) to 16.40% (4B), before narrowing sharply to 3.90% (8B).
Model | |||
InternVL2.5-2B | 70.07 | 40.68 | 55.49 |
InternVL2.5-4B | 75.61 | 45.23 | 61.63 |
InternVL2.5-8B | 74.31 | 57.75 | 61.65 |